BearsGirl:

A delivery truck is transporting boxes of two sizes: large and small. The large boxes weigh 50 pounds each, and the small boxes weigh 35 pounds each. There are 120 boxes in all. If the truck is carrying a total of 5025 pounds in boxes, how many of each type of box is it carrying?

ThisGirlPretty:

@Vocaloid

Vocaloid:

you have to write a system of equations to solve this let L = number of large boxes and S = number of small boxes L + S = 120 since there are 120 boxes and each box must be large or small then we write an equation for total weight (weight of large box)(L) + (weight of small box)(S) = total weight

Vocaloid:

L + S = 120 50L + 35S = 5025 solve by eliminating either L or S

BearsGirl:

thanks